# The Tarantula–Revealed by X-rays (T-ReX)

### v1

@@ Abstract
journal: ApJS — 21 March 2024facilities: CXO (ACIS), Gaia, ESO:VISTA (VIRCAM), VLT:Kueyen (FLAMES-ARGUS), HST (WFC3, ACS), Spitzer (IRAC, MIPS), WISE Herschel (SPIRE)software: ACIS Extract (Broos et al., 2010, 2012; Broos & Townsley, 2016), CIAO (Fruscione et al., 2006), DS9 (Joye & Mandel, 2003), MARX (Davis et al., 2012) XSPEC (Arnaud, 1996).
@@ 2.2.1 Methods, software, and procedures
The algorithms, procedures, and tools used for data reduction, astrometric alignment of observations, point source detection and validation, point source extraction, point source masking, construction of smoothed diffuse emission images, diffuse region extraction, and point source spectral fitting were the same as those used in MOXC2 and MOXC3 (Townsley et al., 2018, 2019). Please see those papers (and their references) for explanations of our data analysis methods. Even though all T-ReX observations have similar aimpoints, we followed our usual practice of extracting point sources and diffuse regions from each observation, and then combining those extractions and calibration products.
@@ 2.3 Point Source Catalog
Extraction apertures for the 3615 point sources in the T-ReX catalog are are overlaid on an image of the full T-ReX event data in Figure 2. The catalog is published as an electronic table accompanying this article, and its columns are summarized in Table 2. The variation in both resolution and sensitivity as a function of increasing off-axis angle $\theta$ are evident in the aperture sizes (increasing) and point-source density (decreasing). Both event data and point-source density decrease further in the S-array detectors farther off-axis where the exposure time is unevenly reduced compared to the field center, due to arbitrary roll angles (Figure 1).
The extraction aperture for each observation of a point source is usually a contour of the local point spread function (PSF), chosen to enclose $\sim$90of the PSF for uncrowded observations or a reduced fraction to handle crowding (Broos et al., 2010). We manually increased aperture sizes for the following very bright sources: VFTS399 (Clark et al., 2015, source Label p1_610), Mk34 (Pollock et al., 2018, p1_995), R140a1/a2 (p1_752), Mk39 (p1_698), R144 (p1_1194), 053844.33-690554.7 (p1_1000), 053844.12-690556.6 (p1_979), SN1987A (c73). Two bright and severely crowded sources, R136a (p1_832) and R136c (p1_893), required hand-drawn apertures (see Fig. 1 of C22, and related discussion of these sources as blends of multiple unresolved massive stars.
@@ 2.7 Archive of Reduced Data Products
The Zenodo archive 10.5281/zenodo.8388113141414https://zenodo.org/record/8388113 (Townsley & Broos, 2023) contains data products and plots that may be of use to other investigators who wish to use the T-ReX data set; see the file README.TXT there for details.
@@ 3 An Imaging Tour of T-ReX
Three neighboring structures, labeled in Figure 2, stand out in the 1.1–2.3 keV band against the softer background. The ${\sim}5$-Myr-old star-forming region NGC 2060 (Sabbi et al., 2016), located ${\sim}6\arcmin$ southwest of R136, is dominated by the young SNR N157B (also known as 30 Dor B), containing the 16-ms PSR J0537-6910, the most energetic pulsar known (Chen et al., 2006, 2023). N157B and its prominent cometary pulsar wind nebula contributed over half of the total X-ray counts in prior Chandra studies of 30 Dor (Townsley et al., 2006a). West of N157B lies the 30 Dor C superbubble, a large shell of non-thermal X-ray emission, which appears as an elongated, broken blue ring surrounding (in projection) a compact, softer source that has been identified as MCSNR J0536–6913 (Kavanagh et al., 2015). A hard-edged artifact produced by smoothing over the mask for the exceptionally bright PSR B0540-69 appears in the S-array mosaic SE of 30 Dor. Enhanced hard diffuse emission in the vicinity is likely associated with the young, oxygen-rich SNR reported by Park et al. (2010) using on-axis Chandra observations.
@@
In these images, the Spitzer/IRAC 8 $\mu$m band (green) shows primarily emission from polycyclic aromatic hydrocarbons (PAHs) that produce a strong 7.7 $\mu$m fluorescence feature when exposed to UV irradiation (Povich et al., 2007), making it an excellent tracer of locations where the surfaces of cold, molecular clouds are exposed to ionizing radiation. The Spitzer/MIPS 24 $\mu$m band (red) traces warm dust, which emits most strongly where dust grains are intermixed with ionized plasma and becomes stochastically heated by the hard radiation field (Watson et al., 2008; Everett & Churchwell, 2010). The full ACIS 0.5–7 keV emission (blue) tends to fill in the cavities and voids in the mid-IR nebula. Regions where the diffuse X-rays are especially faint, notably a broad region bordering the ESE half of 30 Dor (Figure 3), tend to be filled in by MIR cloud structures (Figure 4), indicating that the cold clouds are located in front of the hot plasma in these locations and absorb the soft X-rays. In other regions, diffuse X-rays appear to surround or pass in front of bright MIR filaments (many examples are evident in Figure 5, especially panels c and d), revealing interfaces between the hot and cold ISM.
Refer to captionRefer to captionRefer to captionRefer to caption
@@ 5 Diffuse X-ray Emission
Townsley et al. (2011b, c) provided a detailed description of our general spectral modeling strategy as applied to the Carina Nebula, other Galactic massive star-forming complexes, and the 30 Dor observations predating T-ReX. The spectral model consists of 3 vpshock components; high timescales (e.g. $\log{(\rm tau)}=13.7/[{\rm s~{}cm}^{-3}$]) imply collisional ionization equilibrium (CIE) while lower timescales suggest the presence of non-equilibrium ionization (NEI) plasma. Each spectral fit plot presented below (Figures 911) labels the vpshock parameters for the various employed components: absorbing column (NH[123]), plasma temperature (kT[123]), timescale (tau[123]) and absorption-corrected surface brightness (SB[123]).
Abundances for most elements in the vpshock models were frozen at $0.4Z_{\odot}$, appropriate for the low-metallicity environment of the LMC (Russell & Dopita, 1990), but we found that enhanced abundances of Ne ($0.59Z_{\odot}$), Mg $1.06Z_{\odot}$, and Si ($0.75Z_{\odot}$) were needed to account for other line features in the spectrum. Absorbing columns are modeled as TBabs; the values shown may be doubled to correct for LMC abundances.
@@
We also allow a contribution from unresolved pre-MS stars (2-temperature apec model) and an absorbed power law to account for a hard tail in the spectrum due to a background of unresolved AGN and/or nonthermal emission in 30 Dor.
@@ 5.5 Physical Properties of the Global Diffuse X-ray Emission in 30 Dor
The assumed volume ($\eta V_{x}$) of the hot plasma region in 30 Dor is about 50 times greater than that of Carina, while the total-band, absorption-corrected luminosity ($L_{t,c}$) is ${\sim}120$ times greater (Townsley et al., 2011b). Based on scaling the total energy and mass ($E_{x}$ and $M_{x}$) of the hot plasma in Carina as $(L_{tc}\eta V_{x})^{1/2}$, we would expect these quantities to be roughly 10 times greater in 30 Dor, but instead we find they are 65–75 times greater.
We can also compare the luminosity of the diffuse X-ray emission to the combined bolometric lumnosity of the massive stars in each region. For 30 Dor, $L_{\rm bol}=3\times 10^{41}$ (Lopez et al., 2011) yields $\log{L_{t,c}/L_{\rm bol}}=-4.15$. Adopting $L_{t,c}$ from Townsley et al. (2011c) and $L_{\rm bol}$ from (Binder & Povich, 2018), we have $\log{L_{t,c}/L_{\rm bol}}=-5.70$, --5.52, and --6.04 for Carina, NGC 3603, and M17, respectively, in the Milky Way.
@@ 6 Legacy Value of T–ReX
In this overview of the T-ReX project, we have provided numerous data products, including a point-source catalog, diffuse emission images and spatially-resolved spectra (Section 2.7). The analysis presented in this work only scratches the surface of this unprecedentedly deep and rich X-ray dataset. T-ReX provides our best opportunity to dissect the high-energy emission from a starburst region. These galactic building blocks all emit X-rays, but they blur together in distant starburst galaxies. With T-ReX, we have quantified the relative contributions of the diverse sources that constitute the X-ray emission of a starburst. The stars themselves (single very massive stars, magnetic OB stars, colliding-wind binaries, and lower-mass pre-MS stars) and their stellar remnants (pulsars, X-ray binaries) are generally resolved and hence can be separated from truly diffuse shock emission (from massive stellar winds, individual young SNRs, thermalized cavity SNRs, and PWNe). T-ReX provides the largest number of diffuse X-ray emission counts ever detected from a massive star-forming region, surpassing Carina (Townsley et al., 2011c) and Cygnus OB2 (Albacete-Colombo et al., 2023).
@@
We thank Breanna Binder and Anna Rosen for helpful discussions, offering their perspectives as, respectively, an extragalactic high-energy observer and a theoretician of massive stellar feedback processes. This work was supported by the Chandra X-ray Observatory Guest Observer grants GO5-6080X (PI: L. Townsley) and GO4-15131X (PI: L. Townsley) and by the Penn State ACIS Instrument Team Contract SV4-74018. Both of these were issued by the Chandra X-ray Center, which is operated by the Smithsonian Astrophysical Observatory for and on behalf of NASA under contract NAS8-03060. M. Povich acknowledges support from the National Science Foundation under award AST-2108349 (RUI). The ACIS Guaranteed Time Observations included here were selected by the ACIS Instrument Principal Investigator, Gordon P. Garmire, of the Huntingdon Institute for X-ray Astronomy, LLC, which is under contract to the Smithsonian Astrophysical Observatory; Contract SV2-82024. This research used data products from the Chandra Data Archive, software provided by the Chandra X-ray Center in the application package CIAO, and the SAOImage DS9 software developed by the Smithsonian Astrophysical Observatory. This research also used data products from the Spitzer Space Telescope, operated by the Jet Propulsion Laboratory (California Institute of Technology) under a contract with NASA. This work used data products from the NASA/ESA Hubble Space Telescope, obtained at the Space Telescope Science Institute, which is operated by AURA Inc., under NASA contract NAS 5-26555. This work also used data from the ESA mission Gaia (https://www.cosmos.esa.int/gaia), processed by the Gaia Data Processing and Analysis Consortium (DPAC, https://www.cosmos.esa.int/web/gaia/dpac/consortium); funding for the DPAC has been provided by national institutions, in particular the institutions participating in the Gaia Multilateral Agreement. This research has also made use of NASA’s Astrophysics Data System Bibliographic Services, and the SIMBAD and VizieR databases operated at the Centre de Données Astronomique de Strasbourg. .

### v2

@@ Abstract
journal: ApJSfacilities: CXO (ACIS), Gaia, ESO:VISTA (VIRCAM), VLT:Kueyen (FLAMES-ARGUS), HST (WFC3, ACS), Spitzer (IRAC, MIPS), WISE Herschel (SPIRE)software: ACIS Extract (Broos et al., 2010, 2012; Broos & Townsley, 2016), CIAO (Fruscione et al., 2006), DS9 (Joye & Mandel, 2003), MARX (Davis et al., 2012) XSPEC (Arnaud, 1996).
@@ 2.2.1 Methods, software, and procedures
The algorithms, procedures, and tools used for data reduction, astrometric alignment of observations, point source detection and validation, point source extraction, point source masking, construction of smoothed diffuse emission images, diffuse region extraction, and point source spectral fitting were the same as those used in the second and third installments of our Massive star-forming regions Omnibus X-ray Catalog (MOXC2 and MOXC3, Townsley et al., 2018, 2019). Please see those papers (and their references) for explanations of our data analysis methods. Even though all T-ReX observations have similar aimpoints, we followed our usual practice of extracting point sources and diffuse regions from each observation, and then combining those extractions and calibration products.
@@ 2.3 Point Source Catalog
Extraction apertures for the 3615 point sources in the T-ReX catalog are overlaid on an image of the full T-ReX event data in Figure 2. The catalog is published as an electronic table accompanying this article, and its columns are summarized in Table 2. The variation in both resolution and sensitivity as a function of increasing off-axis angle $\theta$ are evident in the aperture sizes (increasing) and point-source density (decreasing). Both event data and point-source density decrease further in the S-array detectors farther off-axis where the exposure time is unevenly reduced compared to the field center, due to arbitrary roll angles (Figure 1).
The extraction aperture for each observation of a point source is usually a contour of the local point spread function (PSF), chosen to enclose $\sim$90of the PSF for uncrowded observations or a reduced fraction to handle crowding (Broos et al., 2010). We manually increased aperture sizes for the following very bright sources: VFTS399 (Clark et al., 2015, source Label p1_610), Mk34 (Pollock et al., 2018, p1_995), R140a1/a2 (p1_752), Mk39 (p1_698), R144 (p1_1194), 053844.33-690554.7 (p1_1000), 053844.12-690556.6 (p1_979), SN1987A (c73). Two bright and severely crowded sources, R136a (p1_832) and R136c (p1_893), required hand-drawn apertures (see Fig. 1 of C22, and related discussion of these sources as blends of multiple unresolved massive stars).
@@ 2.7 Archive of Reduced Data Products
The Zenodo archive (Townsley & Broos, 2023) [^14] https://zenodo.org/doi/10.5281/zenodo.8372420 contains data products and plots that may be of use to other investigators who wish to use the T-ReX data set; see the file README.TXT there for details.
@@ 3 An Imaging Tour of T-ReX
Refer to captionRefer to captionRefer to captionRefer to caption
Three neighboring structures, labeled in Figure 2, stand out in the 1.1–2.3 keV band against the softer background. The ${\sim}5$-Myr-old star-forming region NGC 2060 (Sabbi et al., 2016), located ${\sim}6\arcmin$ southwest of R136, is dominated by the young SNR N157B (also known as 30 Dor B), containing the 16-ms PSR J0537-6910 (Chen et al., 2006, 2023). N157B and its prominent cometary pulsar wind nebula contributed over half of the total X-ray counts in prior Chandra studies of 30 Dor (Townsley et al., 2006a). West of N157B lies the 30 Dor C superbubble, a large shell of non-thermal X-ray emission, which appears as an elongated, broken blue ring surrounding (in projection) a compact, softer source that has been identified as MCSNR J0536–6913 (Kavanagh et al., 2015). A hard-edged artifact produced by smoothing over the mask for the exceptionally bright PSR B0540-69 appears in the S-array mosaic SE of 30 Dor. Enhanced hard diffuse emission in the vicinity is likely associated with the young, oxygen-rich SNR reported by Park et al. (2010) using on-axis Chandra observations.
@@
In these images, the Spitzer/IRAC 8 $\mu$m band (green) shows primarily emission from polycyclic aromatic hydrocarbons (PAHs) that produce a strong 7.7 $\mu$m fluorescence feature when exposed to UV irradiation (Povich et al., 2007), making it an excellent tracer of locations where the surfaces of cold, molecular clouds are exposed to ionizing radiation. The Spitzer/MIPS 24 $\mu$m band (red) traces warm dust, which emits most strongly where dust grains are intermixed with ionized plasma and becomes stochastically heated by the hard radiation field (Watson et al., 2008; Everett & Churchwell, 2010). The full ACIS 0.5–7 keV emission (blue) tends to fill in the cavities and voids in the mid-IR nebula. Regions where the diffuse X-rays are especially faint, notably a broad region bordering the East-Southeast half of 30 Dor (Figure 3), tend to be filled in by mid-IR cloud structures (Figure 4), indicating that the cold clouds are located in front of the hot plasma in these locations and absorb the soft X-rays. In other regions, diffuse X-rays appear to surround or pass in front of bright mid-IR filaments (many examples are evident in Figure 5, especially panels c and d), revealing interfaces between the hot and cold ISM.
@@ 5 Diffuse X-ray Emission
Townsley et al. (2011b, c) provided a detailed description of our general spectral modeling strategy as applied to the Carina Nebula, other Galactic massive star-forming complexes, and the 30 Dor observations predating T-ReX. The spectral model consists of three shocked thermal plasma components implemented in XSPEC (vpshock). Components 1 and 2 consist of cooler plasmas ($0.25~{}{\rm keV}\lesssim kT\lesssim 0.35$ keV) with different cooling timescales; the high timescales for component 2 (e.g. $\log{(\rm\tau)}=13.7/[{\rm s~{}cm}^{-3}$]) imply collisional ionization equilibrium (CIE) while the lower timescales for component 1 suggest the presence of non-equilibrium ionization (NEI) plasma. Component 3 consists of hotter plasmas ($0.7~{}{\rm keV}\lesssim kT\lesssim 1.3$ keV) exhibiting a mixture of CIE and NEI. Each spectral fit plot presented below (Figures 911) labels the vpshock parameters for the various employed components: absorbing column $N_{H}$ is labeled as NH[123], plasma temperature $kT$ as kT[123], timescale $\tau$ as tau[123] and absorption-corrected surface brightness as SB[123].
We assumed that absorbing material in both the LMC and foreground Milky Way have relative abundances equal to those of the Sun, but that the LMC abundances are scaled down from the Milky Way (by about 50%). Thus in XSPEC we chose to model total absorption with a single TBabs model that uses solar abundances. Abundances for most elements in the vpshock models were frozen at $0.4Z_{\odot}$, appropriate for the low-metallicity environment of the LMC (Russell & Dopita, 1990), but we found that enhanced abundances of Ne ($0.59Z_{\odot}$), Mg $1.06Z_{\odot}$, and Si ($0.75Z_{\odot}$) were needed to account for other line features in the spectrum. The NH[123] values labeled in Figures 911 may be doubled to correct for LMC abundances.
@@
We also allow a contribution from unresolved pre-MS stars (2-temperature apec model for collisionally-ionized thermal plasma) and an absorbed power law to account for a hard tail in the spectrum due to a background of unresolved AGN and/or nonthermal emission in 30 Dor.
@@ 5.5 Physical Properties of the Global Diffuse X-ray Emission in 30 Dor
Because both the three-dimensional geometry and filling factor of the X-ray emitting plasma are unknown, we introduce a correction factor $\eta$ that can range from zero to unity that applies to all physical quantities related to the volume of the hot plasma (see Table 3). The assumed volume ($\eta V_{x}$) of the hot plasma region in 30 Dor is about 50 times greater than that of Carina, while the total-band, absorption-corrected luminosity ($L_{tc}$) is ${\sim}120$ times greater (Townsley et al., 2011b). Both the total energy and mass ($E_{x}$ and $M_{x}$) of the hot plasma scale as $(L_{tc}\eta V_{x})^{1/2}$. From Townsley et al. (2003), $L_{tc}=\Lambda n_{e,x}^{2}\eta V_{x}$, where $\Lambda=5\times 10^{-23}$ erg cm3 s-1 is the bolometric X-ray emissivity (Landi & Landini, 1999). Then the total thermal energy can be written as
$E_{x}=\frac{3}{2}P_{x}\eta V_{x}\propto n_{e,x}T_{x}\eta V_{x}\propto\left(\frac{L_{tc}}{\Lambda\eta V_{x}}\right)^{1/2}\eta V_{x}T_{x}=(L_{tc}\eta V_{x}/\Lambda)^{1/2}T_{x}.$ (2)
The total mass $M_{x}=\mu_{x}m_{p}n_{e}\eta V_{x}$ (denoting the mean particle mass as $\mu_{x}m_{p}$) has the same dependence on number density and emitting volume as the total energy, and hence also scales with luminosity as shown in Equation 2. Based on the above scaling relations and the very similar observed plasma temperatures, we would expect the total energy and mass of X-ray emitting plasam to be roughly 10 times greater in 30 Dor than in Carina, but instead we find they are 65–75 times greater.
We can also compare the luminosity of the diffuse X-ray emission to the combined bolometric lumnosity of the massive stars in each region. For 30 Dor, $L_{\rm bol}=3\times 10^{41}$ (Lopez et al., 2011) yields $\log{L_{tc}/L_{\rm bol}}=-4.15$. Adopting $L_{tc}$ from Townsley et al. (2011c) and $L_{\rm bol}$ from (Binder & Povich, 2018), we have $\log{L_{tc}/L_{\rm bol}}=-5.70$, --5.52, and --6.04 for Carina, NGC 3603, and M17, respectively, in the Milky Way.
@@ 6 Legacy Value of T–ReX
In this overview of the T-ReX project, we have provided numerous data products, including a point-source catalog, diffuse emission images and spatially-resolved spectra (Section 2.7). The analysis presented in this work only scratches the surface of this unprecedentedly deep and rich X-ray dataset. T-ReX provides our best opportunity to dissect the high-energy emission from a starburst region. These galactic building blocks all emit X-rays, but they blur together in distant starburst galaxies. With T-ReX, we have quantified the relative contributions of the diverse sources that constitute the X-ray emission of a starburst. The stars themselves (single very massive stars, colliding-wind binaries, and lower-mass pre-MS stars) and their stellar remnants (pulsars, X-ray binaries) are generally resolved and hence can be separated from truly diffuse shock emission (from massive stellar winds, individual young SNRs, thermalized cavity SNRs, and PWNe). T-ReX provides the largest number of diffuse X-ray emission counts ever detected from a massive star-forming region, surpassing Carina (Townsley et al., 2011c) and Cygnus OB2 (Albacete-Colombo et al., 2023).
@@
We appreciate the time and effort donated by our anonymous referee to provide helpful and supportive comments on this paper. We thank Breanna Binder and Anna Rosen for helpful discussions, offering their perspectives as, respectively, an extragalactic high-energy observer and a theoretician of massive stellar feedback processes. This work was supported by the Chandra X-ray Observatory Guest Observer grants GO5-6080X (PI: L. Townsley) and GO4-15131X (PI: L. Townsley) and by the Penn State ACIS Instrument Team Contract SV4-74018. Both of these were issued by the Chandra X-ray Center, which is operated by the Smithsonian Astrophysical Observatory for and on behalf of NASA under contract NAS8-03060. M. Povich acknowledges support from the National Science Foundation under award AST-2108349 (RUI). The ACIS Guaranteed Time Observations included here were selected by the ACIS Instrument Principal Investigator, Gordon P. Garmire, of the Huntingdon Institute for X-ray Astronomy, LLC, which is under contract to the Smithsonian Astrophysical Observatory; Contract SV2-82024. This research used data products from the Chandra Data Archive, software provided by the Chandra X-ray Center in the application package CIAO, and the SAOImage DS9 software developed by the Smithsonian Astrophysical Observatory. This research also used data products from the Spitzer Space Telescope, operated by the Jet Propulsion Laboratory (California Institute of Technology) under a contract with NASA. This work used data products from the NASA/ESA Hubble Space Telescope, obtained at the Space Telescope Science Institute, which is operated by AURA Inc., under NASA contract NAS 5-26555. This work also used data from the ESA mission Gaia (https://www.cosmos.esa.int/gaia), processed by the Gaia Data Processing and Analysis Consortium (DPAC, https://www.cosmos.esa.int/web/gaia/dpac/consortium); funding for the DPAC has been provided by national institutions, in particular the institutions participating in the Gaia Multilateral Agreement. This research has also made use of NASA’s Astrophysics Data System Bibliographic Services, and the SIMBAD and VizieR databases operated at the Centre de Données Astronomique de Strasbourg. .